OpenStudy (anonymous):

Given f(x)=|x-3|-5, find the average rate of change of the function on [1,5].

terenzreignz (terenzreignz):

You'll need to evaluate the function at the endpoints of that interval first...

terenzreignz (terenzreignz):

--- What are f(1) and f(5) ?

OpenStudy (anonymous):

How do I do that? I'm sorry stupid here:( this is all new to me

terenzreignz (terenzreignz):

You don't know how to evaluate a function at specific points? Here's an example, suppose we have g(x) = 2x + 1 g(0) = 2(0) + 1 [replace the x with 0, then evaluate ] = 0 + 1 = 1 g(1) = 2(1) + 1 [replace the x with 1, then evaluate ] = 2 + 1 = 3 Now, just do the same with f(x) = |x-3|-5

terenzreignz (terenzreignz):

And by the way, once you've done that, the average rate of change of a function between two points \(\large x_1\) and \(\large x_2\) is \[\Large \frac{f(x_2)-f(x_1)}{x_2-x_1}\] It's just like finding the slope of a line, between two points, see? ^_^

OpenStudy (anonymous):

I'll try my best that makes more sense now that you've explained it and given me a rate of change function thanks:)

terenzreignz (terenzreignz):

When you're done, please post your answer and (preferably) how you got it. Thanks ^_^